### v1

@@ 3.3 A path with an odd number of points
In this section, we obtain a noncrossing longest path with an odd number of points. Here is the place where we use the coordinate $(0,0)$ of the point $p_{1}\in P$. We show that our construction for even paths leads to a construction for odd paths by simply removing $p_{1}$. Thus we do not need to repeat the lemmas of Section 3.2 for the odd case.
We claim that if we remove the point $p_{1}$ from the path $H^{\prime}$ constructed on $P^{\prime}$ in the previous section, the remaining path, i.e., $H^{\prime\prime}=p_{-2},p_{2},p_{-3},p_{3},\dots,p_{k},p_{-1}$, is the longest path for the remaining $12k{-}$ 1 points. By construction, $|H^{\prime}|=|H^{\prime\prime}|+|p_{1}p_{-2}|$. Assume, for the sake of proof by contradiction, that the longest path $L$ for the remaining points is longer than $H^{\prime\prime}$. Among the two endpoints of $L$, let $p_{i}$ be an endpoint that is not $p_{-1}$. Due to our choices of the $x$- and $y$-coordinates we have $|p_{1}p_{i}|\geq|p_{1}p_{-2}|$. Therefore the concatenation of $L$ and $p_{1}p_{i}$ would give a path on $P^{\prime}$ of length $|L|+|p_{1}p_{-2}|$ which is larger than $|H^{\prime}|$. This contradicts $H^{\prime}$ being the longest path on $P^{\prime}$.
@@ 4.1 A cycle with an even number of points: An overview
Let $n\geq 6$ be an even number. Then either $n=4k$ or $n=4k{-}2$ for some integer $k$. To simplify the indexing (of points and $y$-coordinates) in our construction, from now on we assume that $n=4k{-}2$. Let $P$ be a set of $n$ points, consisting of $2k$ points $p_{i}=(i,0)$ for $i=\pm 1,\pm 2,\dots,\pm k$ and $22k{-}$ 2 points $p^{\prime}_{i}=(i{+}\epsilon,0)$ for $i=-1,\pm 2,\dots,\pm(k{-}1),k$, where $\epsilon>0$ is a small value to be determined; see Figure 3. (The construction for $n=4k$ is similar; it consists of $P$ and two additional points $p_{k+1}=(k{+}1,0)$ and $p^{\prime}_{-k}=(-k{+}\epsilon,0)$.) Our construction for cycles is somewhat similar to that of paths in the sense that our cycle consists of two $y$-monotone interior-disjoint paths between $p_{1}$ and $p_{-k}$ (or between $p_{1}$ and $p_{k+1}$ when $n$ is a multiple of $4$). Although the main idea sounds simple, the noncrossing property of the longest cycle is not straightforward and involves a more detailed analysis.
@@
Due to the small $y$-coordinates, any longest cycle $C^{\prime}$ on $P^{\prime}$ corresponds to a cycle $C\in\cal C$. Moreover $|C^{\prime}|=|C|+\Delta(C^{\prime})$ for some small value $\Delta(C^{\prime})$ which depends on the new $y$-coordinates. Since $p_{1}$ has the largest $y$-coordinate, the contribution of the two edges of $C^{\prime}$ that are incident to $p_{1}$ (say $e_{1}$ and $e_{2}$) is maximized when they are connected to the nearest plausible neighbors which are $p_{-1}$ and $p^{\prime}_{-1}$. We will choose the $y$-coordinates in such a way that the contribution of $e_{1}$ and $e_{2}$ is larger than the sum of the contributions of the remaining edges of the cycle. Thus $C^{\prime}$ must connect $p_{1}$ to $p_{-1}$ and $p^{\prime}_{-1}$. Similarly, by a suitable choice of $y$-coordinates, we enforce $C^{\prime}$ to connect $p_{-1}$ and $p^{\prime}_{-1}$ to the nearest plausible neighbors which are $p_{2}$ and $p^{\prime}_{2}$, and so on. By repeating this process, the longest cycle $C^{\prime}$ would be the concatenation of two paths $p_{1},p_{-1},p_{2},p_{-2},\dots,p_{-k}$ and $p_{1},p^{\prime}_{-1},p^{\prime}_{2},p^{\prime}_{-2},\dots,p^{\prime}_{k},p_{-k}$.
Figure 3: Illustration of the construction of a longest cycle for $24k{-}$ 2 points. The figure is not to scale. The $y$-coordinates should be small enough so that all points lie almost on the $x$-axis.
@@ 4.2 A cycle with an even number of points: Details
Recall the point set $P$ from the previous section (the $y$-coordinates and the value of $\epsilon>0$ will be determined in this section). The longest cycles for points on a line were characterized in Lemma 3. Let $\cal C$ be the set of all longest cycles on $P$.
@@
Any cycle in $\cal C$ is at least 1 unit longer than any cycle not in $\cal C$.
@@
Consider any cycle $D$ that is not in $\cal C$. Lemma 3 implies that $D$ has an edge that does not intersect the $y$-axis. Orient the edges of $D$ to make it a directed cycle. Since the number of points to the left of the $y$-axis is the same as the number of points to its right, $D$ has two directed edges $(p_{a},p_{b})$ and $(p_{c},p_{d})$ such that $a,b\leq-1$ and $c,d\geq 1$. By replacing these edges with $p_{a}p_{c}$ and $p_{b}p_{d}$ we obtain an (undirected) spanning cycle $D^{\prime}$ such that
$|D^{\prime}|-|D|=(|p_{a}p_{c}|+|p_{b}p_{d}|)-(|p_{a}p_{b}|+|p_{c}p_{d}|)\geq 2|p_{1}p^{\prime}_{-1}|=2(2-\epsilon)>1.$
Since the length of any cycle $C$ in $\cal C$ is at least $|D^{\prime}|$, we get $|C|>|D|+1$. ∎
Lemma 7.
Let $0\leq\varepsilon\leq 1/16k$ be a real number. Suppose that every point of $P$ is perturbed by a distance of at most $\varepsilon$. Then the order of the points along any longest cycle of the new point set is the same as the order of the points along some cycle in $\cal C$.
Proof Sketch.
The proof is similar to that of Lemma 4 and uses Lemma 6. The parameter $\varepsilon$ is small enough such that the total change in the length of any spanning cycle on $P$ is less than 1/2. Together with Lemma 6, this implies that any longest cycle on the perturbed points corresponds to a cycle in $\cal C$.
@@
The $y$-coordinates of the remaining points would then follow as outlined in the previous section (more details are given after Lemma 9). We set $y_{1}=\frac{1}{16k}$ and $y_{-k}=0$. We use the following lemma (which can be proven similarly to Lemma 5) to assign the $y$-coordinates.
@@
Lemma 7 implies that any longest cycle $C^{\prime}$ on $P^{\prime}$ corresponds to a cycle $C$ in $\cal C$. Due to small $y$-coordinates, we have $|C^{\prime}|=|C|+\Delta(C^{\prime})$ for some small value $\Delta(C^{\prime})\geq 0$ which depends on the new $y$-coordinates. Lemma 3 implies that $C^{\prime}$ connects $p_{1}$ to two points to the left of the $y$-axis. Similar to Lemma 5, for $j\in\{1,\dots,k\}$ define $\ell_{j}$ as the Euclidean distance between $p_{1}$ and the point $(-j,0)$, and define $\Delta_{j}$ as the difference of their $x$-coordinates. Analogously, for $j\in\{1,\dots,k{-}1\}$ define $\ell^{\prime}_{j}$ and $\Delta^{\prime}_{j}$ for $p_{1}$ and the point $(0,-j+\epsilon)$. Every edge that connects $p_{1}$ to a point to the left of the $y$-axis has the following contributions to $|C|$, $|C^{\prime}|$ and $\Delta(C^{\prime})$.
  • For $j\in\{1,\dots,k\}$ the contribution of $p_{1}p_{-j}$ to $|C^{\prime}|$ is at least $\ell_{j}-\delta$ and at most $\ell_{j}$. The contribution of the corresponding edge to $|C|$ is $\Delta_{j}$. Hence the contribution of $p_{1}p_{-j}$ to $\Delta(C^{\prime})$ is at least $\ell_{j}-\delta-\Delta_{j}$ and at most $\ell_{j}-\Delta_{j}$.
  • For $j\in\{2,\dots,k{-}1\}$ the contribution of $p_{1}p^{\prime}_{-j}$ to $|C^{\prime}|$ is at least $\ell^{\prime}_{j}-\delta$ and at most $\ell_{j}$. The contribution of the corresponding edge to $|C|$ is $\Delta^{\prime}_{j}$. Thus the contribution of $p_{1}p^{\prime}_{-j}$ to $\Delta(C^{\prime})$ is at least $\ell^{\prime}_{j}-\delta-\Delta^{\prime}_{j}$ and at most $\ell^{\prime}_{j}-\Delta^{\prime}_{j}$.
  • The contribution of $p_{1}p^{\prime}_{-1}$ to $|C^{\prime}|$ is at least $\ell^{\prime}_{1}-\delta-\epsilon$ because the $y$-coordinate of $p^{\prime}_{-1}$ is at most $\delta+\epsilon$; to verify this observe that $y_{-1}\leq\delta$ and $y^{\prime}_{-1}-y_{-1}<\epsilon$ because $p^{\prime}_{-1}$ is almost on $p_{1}p_{-1}$ whose slope is less than $1$; also see Figure 3 (recall that the figure is not to scale). The contribution of the corresponding edge to $|C|$ is $\Delta^{\prime}_{1}$. Therefore the contribution of $p_{1}p^{\prime}_{-1}$ to $\Delta(C^{\prime})$ is at least $\ell^{\prime}_{1}-\delta-\epsilon-\Delta^{\prime}_{1}$ and at most $\ell^{\prime}_{1}-\Delta^{\prime}_{1}$.
@@
If we set $\delta<\frac{1}{2}\left((\ell_{1}-\Delta_{1})-(\ell^{\prime}_{2}-\Delta^{\prime}_{2})\right)$, then the contributions of $p_{1}p_{-1}$ and $p_{1}p^{\prime}_{-1}$ to $\Delta(C^{\prime})$ would respectively be at least
@@
which are larger than the contribution of any other edge $p_{1}p_{-j}$ and $p_{1}p^{\prime}_{-j}$. By setting
$\delta=\frac{1}{2}\frac{(\ell_{1}-\Delta_{1})-(\ell^{\prime}_{2}-\Delta^{\prime}_{2})}{4k-2}$
the contribution of each of $p_{1}p_{-1}$ and $p_{1}p^{\prime}_{-1}$ would be even larger than the sum of the contributions of the remaining $44k{-}$ 4 edges of $C^{\prime}$. Thus, for this choice of $\delta$ the longest cycle $C^{\prime}$ connects $p_{1}$ to $p_{-1}$ and $p^{\prime}_{-1}$. ∎
@@
There exists a real $\delta$, $\epsilon\leq\delta<y_{-1}$, such that if $0\leq y_{i}\leq\delta$ for $i\neq-1,1,2$ and $0\leq y^{\prime}_{i}\leq\delta$ for $i\neq-1$, then every longest cycle of $P^{\prime}$ connects $p_{-1}$ to $p_{2}$ and $p^{\prime}_{-1}$ to $p^{\prime}_{2}$.
@@
Recall the longest cycle $C^{\prime}$ from the proof of Lemma 8. We choose $\delta$ small enough such that the contribution of each of $p_{-1}p_{2}$, $p_{-1}p^{\prime}_{2}$, $p^{\prime}_{-1}p_{2}$, and $p^{\prime}_{-1}p^{\prime}_{2}$ to $\Delta(C^{\prime})$ is larger than the sum of the contributions of the remaining $64k{-}$ 6 edges of $C^{\prime}$. This would force $C^{\prime}$ to connect $p_{-1}$ and $p^{\prime}_{-1}$ to $p_{2}$ and $p^{\prime}_{2}$.
@@
We choose our new $\delta$ as in the proof of Lemma 9, and set $y^{\prime}_{2}=\delta$. Now that the point $p^{\prime}_{2}$ is fixed we can choose the $y$-coordinate of $p_{2}$ in the triangle $\bigtriangleup p_{-1}p^{\prime}_{-1}p^{\prime}_{2}$ and very close to the segment $p^{\prime}_{-1}p^{\prime}_{2}$ such that (1) holds. This forces the longest cycle to use $p_{-1}p_{2}$ and $p^{\prime}_{-1}p^{\prime}_{2}$. By repeatedly applying Lemma 9, the longest cycle will use the edges $p_{i}p_{-i}$ and $p^{\prime}_{i}p^{\prime}_{-i}$ (for $i>0$) and the edges $p_{i}p_{-i+1}$ and $p^{\prime}_{i}p^{\prime}_{-i+1}$ (for negative $i<0$). Therefore the longest cycle on $P^{\prime}$ is the concatenation of two paths: $p_{1},p_{-1},p_{2},p_{-2},\dots,p_{-k}$ and $p_{1},p^{\prime}_{-1},p^{\prime}_{2},p^{\prime}_{-2},\dots,p^{\prime}_{k},p_{-k}$. This cycle is unique and noncrossing.
Each time we apply Lemma 9 we obtain a new value for $\delta$. In each application we need $\delta$ to be greater than or equal to our fixed parameter $\epsilon$. For this purpose, we choose $\epsilon$ to be the parameter $\delta$ that is obtained in the last application of Lemma 9, i.e., $\delta=y^{\prime}_{k}$.

### v2

@@ 3.2 A Path with an Odd Number of Points
In this subsection, we extend Lemma 6 to odd point sets, by reducing it to the even case.
Lemma 9.
For an odd integer $n\geq 3$, let $P$ be a set of $n$ real numbers such that the median of $P$ has a unique closest point in $P$. Then there exists a set $P^{\prime}$ of $n$ points in the plane with the following properties:
  1. 1.
    the $x$-projection of $P^{\prime}$ is $P$;
  2. 2.
    the $x$-projection of every longest path on $P^{\prime}$ is a longest path on $P$;
  3. 3.
    the longest spanning path on $P^{\prime}$ is unique and noncrossing.
Proof.
Assume that n=2k1n=2k-1 for some $k\in\mathbb{N}$ and $P=\{x_{-k},x_{-(k-1)}\ldots,x_{k}\}$, where $i<j$ implies $x_{i}<x_{j}$. Note that $x_{0}$ is the median of $P$. We may assume without loss of generality that $x_{0}=0$, and $|x_{-1}|<|x_{1}|$ by symmetry. Let $x_{1/2}=\frac{1}{2}\cdot x_{1}$ and $Q=P\cup\{x_{1/2}\}$. Note that $Q$ contains an even number of points and the closest points to its median are $x_{0}$ and $x_{1/2}$. By Lemma 6, we can assign $y$-coordinates to the points in $Q$ to obtain a point set $Q^{\prime}$ for which the longest spanning path $H^{\prime}$ is unique, $y$-monotone, and its endpoints are $p_{0}$ and $p_{1/2}=(x_{1/2},y_{1/2})$. By the choice of $y$-coordinates, the longest path on every subset of $Q^{\prime}$ corresponds to a longest path on $Q$. We may further assume that $p_{1/2}$ has the maximum $y$-coordinate in $Q^{\prime}$, and $p_{1/2}p_{-1}$ is an edge of $H^{\prime}$.
Let $P^{\prime}=Q^{\prime}\setminus\{p_{1/2}\}$, the subset of $Q^{\prime}$ corresponding to $P$. Since $x_{1/2}$ is an endpoint of $H^{\prime}$, if we remove $x_{1/2}$ from $H^{\prime}$, we obtain a spanning path $H^{\prime\prime}$ on $P^{\prime}$, with endpoints $p_{0}$ and $p_{-1}$. We claim that $H^{\prime\prime}$ is a longest spanning path on $P^{\prime}$. Suppose, for the sake of contradiction, that $P^{\prime}$ admits a path $L$ that is longer than $H^{\prime\prime}$. By our choice of $y$-coordinates, the $x$-projection of both $H^{\prime\prime}$ and $L$ is a longest path on $P$. By Lemma 4, the endpoints of both $H^{\prime\prime}$ and $L$ are $p_{-1}$ and $p_{0}$. Consequently, the concatenation of $L$ and the edge $p_{0}p_{1/2}$ gives a spanning path on $Q^{\prime}$: its length is $|L|+|p_{0}p_{1/2}|>|H^{\prime}|+|p_{0}p_{1/2}|=|H^{\prime}|$, contradicting the maximality of $H^{\prime}$. ∎
@@ 4.1 A Cycle with an Even Number of Points: An Overview
Let $n\geq 6$ be an even integer. Then either $n=4k$ or n=4k2n=4k{-}2 for some integer $k\geq 2$. To simplify the indexing (of points and $y$-coordinates) in our construction, from now on we assume that n=4k2n=4k{-}2. Let $P$ be a set of $n$ points, consisting of $2k$ points $p_{i}=(i,0)$ for $i=\pm 1,\pm 2,\dots,\pm k$ and 2k22k{-}2 points $p^{\prime}_{i}=(i{+}\epsilon,0)$ for $i=-1,\pm 2,\dots,\pm(k{-}1),k$, where $\epsilon>0$ is a small value to be determined; see Figure 3. (The construction for $n=4k$ is similar; it consists of $P$ and two additional points $p_{k+1}=(k{+}1,0)$ and $p^{\prime}_{-k}=(-k{+}\epsilon,0)$.) Our construction for cycles is somewhat similar to that of paths in the sense that our cycle consists of two $y$-monotone interior-disjoint paths between $p_{1}$ and $p_{-k}$ (or between $p_{1}$ and $p_{k+1}$ when $n$ is a multiple of $4$). Although the main idea sounds simple, the noncrossing property of the longest cycle is not straightforward and involves a more detailed analysis. The construction could also be done with more general $x$-coordinates as for the paths, but for simplicity we pick (almost) integers.
@@
Due to the small $y$-coordinates, every longest cycle $C^{\prime}$ on $P^{\prime}$ corresponds to a cycle $C\in\cal C$. Moreover $|C^{\prime}|=|C|+Y(C^{\prime})$ for some small value $Y(C^{\prime})$ which depends on the new $y$-coordinates. Since $p_{1}$ has the largest $y$-coordinate, the contribution of the two edges of $C^{\prime}$ that are incident to $p_{1}$ (say $e_{1}$ and $e_{2}$) is maximized when they are connected to the nearest plausible neighbors which are $p_{-1}$ and $p^{\prime}_{-1}$. We will choose the $y$-coordinates in such a way that the contribution of $e_{1}$ and $e_{2}$ is larger than the sum of the contributions of the remaining edges of the cycle. Thus $C^{\prime}$ must connect $p_{1}$ to $p_{-1}$ and $p^{\prime}_{-1}$. Similarly, by a suitable choice of $y$-coordinates, we enforce $C^{\prime}$ to connect $p_{-1}$ and $p^{\prime}_{-1}$ to the nearest plausible neighbors which are $p_{2}$ and $p^{\prime}_{2}$, and so on. By repeating this process, the longest cycle $C^{\prime}$ will be the concatenation of two paths $p_{1},p_{-1},p_{2},p_{-2},\dots,p_{-k}$ and $p_{1},p^{\prime}_{-1},p^{\prime}_{2},p^{\prime}_{-2},\dots,p^{\prime}_{k},p_{-k}$.
Figure 3: Illustration of the construction of a longest cycle for 4k24k{-}2 points. The figure is not true to scale. The $y$-coordinates should be small enough so that all points lie almost on the $x$-axis.
@@ 4.2 A Cycle with an Even Number of Points: Details
Recall the point set $P$ from the previous subsection (the $y$-coordinates and the value of $\epsilon>0$ will be determined in this section). The longest cycles for points on a line were characterized in Lemma 5. Let $\cal C$ be the set of all longest cycles on $P$.
@@
Every cycle in $\cal C$ is more than 2 units longer than every cycle not in $\cal C$.
@@
Consider any cycle $D$ that is not in $\cal C$. Lemma 5 implies that $D$ has an edge that does not intersect the $y$-axis. Orient the edges of $D$ to make it a directed cycle. Since the number of points to the left of the $y$-axis is the same as the number of points to its right, $D$ has two directed edges $(p_{a},p_{b})$ and $(p_{c},p_{d})$ such that $a,b\leq-1+\epsilon$ and $c,d\geq 1$. By replacing these edges with $p_{a}p_{c}$ and $p_{b}p_{d}$ we obtain an (undirected) spanning cycle $D^{\prime}$ such that
$|D^{\prime}|-|D|=(|p_{a}p_{c}|+|p_{b}p_{d}|)-(|p_{a}p_{b}|+|p_{c}p_{d}|)\geq 2|p_{1}p^{\prime}_{-1}|=2(2-\epsilon)>2.$
Since the length of every cycle $C$ in $\cal C$ is at least $|D^{\prime}|$, we get $|C|>|D|+2$. ∎
@@
The $y$-coordinates of the remaining points will then follow as outlined in the previous section (more details are given after Lemma 12). We set $y_{1}=\frac{1}{16k}$ and $y_{-k}=0$. To assign the $y$-coordinates we use the following lemma (its proof is similar to that of Lemma 7). We cannot use Lemma 7 and Lemma 8 directly because the two subpaths forming the cycle are not necessarily longest paths on their respective point sets (mainly because here $p_{-k}$ is an endpoint of the paths as opposed to $p_{-1}$).
@@
Corollary 1 implies that every longest cycle $C^{\prime}$ on $P^{\prime}$ corresponds to a cycle $C$ in $\cal C$. By the triangle inequality and our choice of the $y$-coordinates, we have $|C^{\prime}|=|C|+Y(C^{\prime})$ for $0\leq Y(C^{\prime})\leq|P|\cdot\frac{1}{16k}\leq\frac{1}{4}$. Lemma 5 implies that $C^{\prime}$ connects $p_{1}$ to two points to the left of the $y$-axis. Similar to Lemma 7, for $j\in\{1,\dots,k\}$ define $\ell_{j}$ as the Euclidean distance between $p_{1}$ and the point (j,0)(-j,0), and define $\Delta_{j}$ as the difference of their $x$-coordinates. Analogously, for $j\in\{1,\dots,k{-}1\}$ define $\ell^{\prime}_{j}$ and $\Delta^{\prime}_{j}$ for $p_{1}$ and the point $(0,-j+\epsilon)$. Every edge that connects $p_{1}$ to a point to the left of the $y$-axis has the following contributions to $|C|$, $|C^{\prime}|$ and $Y(C^{\prime})$.
  • For $j\in\{1,\dots,k\}$ the contribution of $p_{1}p_{-j}$ to $|C^{\prime}|$ is at least $\ell_{j}-\delta$ and at most $\ell_{j}$. The contribution of the corresponding edge to $|C|$ is $\Delta_{j}$. Hence the contribution of $p_{1}p_{-j}$ to $Y(C^{\prime})$ is at least $\ell_{j}-\delta-\Delta_{j}$ and at most $\ell_{j}-\Delta_{j}$.
  • For $j\in\{2,\dots,k{-}1\}$ the contribution of $p_{1}p^{\prime}_{-j}$ to $|C^{\prime}|$ is at least $\ell^{\prime}_{j}-\delta$ and at most $\ell_{j}$. The contribution of the corresponding edge to $|C|$ is $\Delta^{\prime}_{j}$. Thus the contribution of $p_{1}p^{\prime}_{-j}$ to $Y(C^{\prime})$ is at least $\ell^{\prime}_{j}-\delta-\Delta^{\prime}_{j}$ and at most $\ell^{\prime}_{j}-\Delta^{\prime}_{j}$.
  • The contribution of $p_{1}p^{\prime}_{-1}$ to $|C^{\prime}|$ is at least $\ell^{\prime}_{1}-\delta-\epsilon$ because the $y$-coordinate of $p^{\prime}_{-1}$ is at most $\delta+\epsilon$; to verify this observe that $y_{-1}\leq\delta$ and $y^{\prime}_{-1}-y_{-1}<\epsilon$ because $p^{\prime}_{-1}$ is almost on $p_{1}p_{-1}$ whose slope is less than $1$; also see Figure 3 (recall that the figure is not to true scale). The contribution of the corresponding edge to $|C|$ is $\Delta^{\prime}_{1}$. Therefore the contribution of $p_{1}p^{\prime}_{-1}$ to $Y(C^{\prime})$ is at least $\ell^{\prime}_{1}-\delta-\epsilon-\Delta^{\prime}_{1}$ and at most $\ell^{\prime}_{1}-\Delta^{\prime}_{1}$.
@@
If we set $\delta<\frac{1}{2}\left((\ell_{1}-\Delta_{1})-(\ell^{\prime}_{2}-\Delta^{\prime}_{2})\right)$, then the contributions of $p_{1}p_{-1}$ and $p_{1}p^{\prime}_{-1}$ to $Y(C^{\prime})$ would respectively be at least
@@
which are larger than the contribution of every other edge $p_{1}p_{-j}$ and $p_{1}p^{\prime}_{-j}$. By setting
$\delta=\frac{1}{2}\cdot\min\left\{\frac{(\ell_{1}-\delta-\Delta_{1})-(\ell^{\prime}_{2}-\Delta^{\prime}_{2})}{4k-2},\frac{(\ell^{\prime}_{1}-\delta-\epsilon-\Delta^{\prime}_{1})-(\ell^{\prime}_{2}-\Delta^{\prime}_{2})}{4k-2}\right\},$
the contribution of $p_{1}p_{-1}$ and $p_{1}p^{\prime}_{-1}$ each would be even larger than the sum of the contributions of the remaining 4k44k{-}4 edges of $C^{\prime}$. Since $\epsilon\leq\delta$, after rearranging we finally fix
$\delta=\min\left\{\frac{(\ell_{1}-\Delta_{1})-(\ell^{\prime}_{2}-\Delta^{\prime}_{2})}{8k-3},\frac{(\ell^{\prime}_{1}-\Delta^{\prime}_{1})-(\ell^{\prime}_{2}-\Delta^{\prime}_{2})}{8k-2}\right\}.$
Thus, for this choice of $\delta$, the longest cycle $C^{\prime}$ connects $p_{1}$ to $p_{-1}$ and $p^{\prime}_{-1}$. ∎
@@
There exists a real number $\delta$, $\epsilon\leq\delta<y_{-1}$, such that if $0\leq y_{i}\leq\delta$ for $i\neq-1,1,2$ and $0\leq y^{\prime}_{i}\leq\delta$ for $i\neq-1$, then every longest cycle of $P^{\prime}$ connects $p_{-1}$ to $p_{2}$ and $p^{\prime}_{-1}$ to $p^{\prime}_{2}$.
@@
Recall the longest cycle $C^{\prime}$ from the proof of Lemma 11. We will choose $\delta$ small enough such that the contribution of each of $p_{-1}p_{2}$, $p_{-1}p^{\prime}_{2}$, $p^{\prime}_{-1}p_{2}$, and $p^{\prime}_{-1}p^{\prime}_{2}$ to $Y(C^{\prime})$ is larger than the sum of the contributions of the remaining 4k64k{-}6 edges of $C^{\prime}$. This will force $C^{\prime}$ to connect $p_{-1}$ and $p^{\prime}_{-1}$ to $p_{2}$ and $p^{\prime}_{2}$.
@@
We choose our new $\delta$ as in the proof of Lemma 12, and set $y^{\prime}_{2}=\delta$. Now that the point $p^{\prime}_{2}$ is fixed we can choose the $y$-coordinate of $p_{2}$ in the triangle $\bigtriangleup p_{-1}p^{\prime}_{-1}p^{\prime}_{2}$ and very close to the segment $p^{\prime}_{-1}p^{\prime}_{2}$ such that (1) holds. This forces the longest cycle to use $p_{-1}p_{2}$ and $p^{\prime}_{-1}p^{\prime}_{2}$. We can repeat the argument of Lemma 12 such that for $j\geq 2$ the longest cycle will use the edges $p_{j}p_{-j}$, $p^{\prime}_{j}p^{\prime}_{-j}$, $p_{-j}p_{j+1}$, and $p^{\prime}_{-j}p^{\prime}_{j+1}$. Therefore the longest cycle on $P^{\prime}$ is the concatenation of two paths: $p_{1},p_{-1},p_{2},p_{-2},\dots,p_{-k}$ and $p_{1},p^{\prime}_{-1},p^{\prime}_{2},p^{\prime}_{-2},\dots,p^{\prime}_{k},p_{-k}$. This cycle is unique and noncrossing.
Each time we apply Lemma 12 we obtain a new value for $\delta$. In each iteration we need $\delta$ to be greater than or equal to our fixed parameter $\epsilon$. For this purpose, we choose $\epsilon$ to be the parameter $\delta$ that is obtained in the last invocation of Lemma 12, that is , $\epsilon=y^{\prime}_{k}$.